Exercise 2.6.22 Given vectors w and x in R, denote their dot product by w x. a. Given w in R, define Tw : R -> R by Tw(x)
Answer #1

2.6.22.

a. TW : Rn→R is defined by TW (x) = w.x for all x in Rn.

Let x and y be 2 arbitrary vectors in Rn and let k be an arbitrary real scalar. Then TW (x+y) = w.(x+y) = w.x+w.y = TW(x)+ TW(y). Thus, TW preserves vector addition.

Also, TW(kx) = w.(kx) = k(w.x) = k TW(x). Thus, TW preserves scalar multiplication.

Hence , TW is a linear transformation.

(b). Let x = (x1,x1,…,xn) and let T Rn→R be a linear transformation defined by T (x) = a1 x1+a2 x2+…+an xn, where a1, a2,…,an are real scalars. Then T(x) = TW (x) where w= (a1, a2,…,an).